OpenStudy (anonymous):

The volume of 10 drops a liquid is 0.1 fluid ounce. What is the volume of 10,000 drops?

hero (hero):

\[\frac{10 \text{drops}}{10,000 \text{drops}} = \frac{0.1 \text{fluid ounces}}{x \space \text{fluid ounces}}\]

hero (hero):

\[\frac{10}{10,000} = \frac{0.1}{x}\] Solve for x